# Cooperative Multi-Agent Planning with Adaptive Skill Synthesis

### v1

@@ 5.2 Main Results
Moreover, COMPASS demonstrates particular advantages in scenarios with sparse reward, where traditional MARL approaches significantly underperform. As shown in Table 3, under sparse reward settings, baseline methods struggle to achieve any meaningful success.
Skill Analysis We now analyze COMPASS’s capability to synthesize and execute diverse tactical behaviors. COMPASS develops four key tactical patterns: (1) An exponentially-scaled focus fire implementation that coordinates multiple units’ target selection based on allied attacker density (Figure 4), (2) A position-aware kiting mechanism that maintains optimal engagement ranges while managing unit positioning relative to threats (Figure 5), (3) A formation-based isolation tactic that enables systematic target elimination through coordinated unit movements (Figure 6), and (4) An area-of-effect (AOE) tactic that maximizes splash damage through cluster density calculation (Figure 7). These synthesized skills exhibit clear strategic intent while maintaining interpretability.
5.3 Ablation Studies
PROTOSS TERRAN ZERG
5V5 $0.35_{0.06}$ $0.24_{0.04}$ $0.06_{0.01}$
5V6 $0.04_{0.05}$ $0.06_{0.02}$ $0.02_{0.03}$
The Effectiveness of Skill Initialization To evaluate the impact of our skill initialization, we analyze the performance of COMPASS using only the initialized skill library derived from expert demonstrations. The results in Table 4 demonstrate that skill initialization alone achieves non-trivial performance across different scenarios, particularly in symmetric matchups. Moreover, the gap between initialized skills and COMPASS underscores the necessity of incremental skill synthesis. A script example for skill initialization is in Appendix.
6 Conclusion
Impact Statement
We believe that the proposed work enhances the capacity for intelligent decision-making in complex and dynamic environments, and can have a positive impact on real-world multi-agent applications such as robotics, traffic management, and resource allocation. However, it is essential to consider potential concerns such as the discrepancy between the simulated environment and the real world. Another potential effect of directly implementing the derived policy is that it could lead to biased decision-making and privacy infringements. Mitigation strategies to address potential hazards could include the establishment of ethical guidelines and regulatory frameworks alongside the integration of transparency.
Appendix A Prompts used in COMPASS
@@ 5. Are there any ally units visible, either observation or minimap?
1. If your chosen region of interest is a unit, format the output as ”[Enemy/Ally] #[target_id]” (e.g., ”Enemy #0” for enemy unit with ID 0, ”Ally #1” for ally unit with ID 1)
2. If your chosen region of interest is location, format the output as ”Location: [direction]” where direction must be one of: ”North”, ”Northeast”, ”East”, ”Southeast”, ”South”, ”Southwest”, ”West”, ”Northwest”, ”Center” (e.g., ”Location: Northeast”)
@@
Prompt for Task Reasoning You are an AI assistant helping with academic research in the StarCraft II’s SMAC (StarCraft Multi-Agent Challenge) environment, controlling a $<$unit_type$>$ unit with ID ¡$unit_{i}d$¿ in micromanagement scenarios $<$scenario_name$>$ to help your team defeat the enemy forces. You operate under decentralized execution with partial observability, making decisions based only on local observations within your unit’s field of view. You will be sequentially given $<$event_count$>$ screenshots and corresponding descriptions of recent events. You will also be given a summary of the history that happened before the last screenshot. By analyzing these inputs, you gain a comprehensive understanding of the current context and situation within the game. You should assist in summarizing the next immediate task to do in SMACv2. Your ultimate goal is to help your team defeat the enemy forces as quickly as possible.
@@
Based on the comprehensive game state analysis and team context, decompose the primary objective of ”defeat all enemy units” into ONE specific tactical sub-task that enhances either target prioritization (score_target) or behavior control (control_logic). This sub-task should be concrete, implementable, and aligned with team coordination. Consider the following in your task decomposition:
@@ 2. Team Context:
”Adjust [scoring weight/multiplier/threshold] to [specific combat calculation] based on [unit composition + battle state] where [precise condition]”
@@
”Implement [unit movement pattern/formation/targeting] when [combat state + ally positions] satisfy [precise conditions]” Task Requirements:
@@
Prompt for Skill Generation You are an AI assistant helping with academic research in the StarCraft II’s SMAC (StarCraft Multi-Agent Challenge) environment, controlling a $<$unit_type$>$ unit with ID ¡$unit_{i}d$¿ in micromanagement scenarios $<$scenario_name$>$ to help your team defeat the enemy forces. You operate under decentralized execution with partial observability, making decisions based only on local observations within your unit’s field of view. Your task is to enhance combat effectiveness:
@@ 1. Analyze the provided script’s effectiveness
2. If the improved function is score_target(unit), there is exactly one parameter named ”unit”.
@@
Prompt for Actor You are an AI assistant helping with academic research in the StarCraft II’s SMAC (StarCraft Multi-Agent Challenge) environment, controlling a $<$unit_type$>$ unit with ID ¡$unit_{i}d$¿ in micromanagement scenarios $<$scenario_name$>$ to help your team defeat the enemy forces. You operate under decentralized execution with partial observability, making decisions based only on local observations within your unit’s field of view. Utilizing this insight, you are tasked with identifying the most suitable skill to take next, given the current task. You control the game unit and can execute skills from the available skill set. Upon evaluating the provided information, your role is to articulate the precise skill you would deploy, considering the game’s present circumstances, and specify any necessary parameters for implementing that skill:
@@
3. The ONLY required keyword parameter is ”obs: str” - you MUST include this parameter as ”obs=’current’” in every skill. The actual observation will be automatically injected at runtime.
@@
def race_melee_ranged_medivac_navi_A_star_score_type_default_center(obs: str):
@@
Zealot/Zergling/Baneling/Colossus/Stalker/Hydralisk/Marauder/Marine/Medivac Controls Logic:
Medivac:
- Heals allies below 100% HP
- Maintains 0.75 sight range from enemies
- Centers between allies when no healing targets
Melee (Zealot/Zergling/Baneling):
- Attacks highest threat target within 0.7 sight range
- Pursues targets using A* pathfinding
- Groups with allies at >0.7 distance threshold
Ranged (Colossus/Stalker/Hydralisk/Marauder/Marine):
- Kites melee units at max_range - 0.05
- Focus fires targets shared with 2+ allies
- Maintains position behind melee allies
Key Implementation:
- Pathfinding: A* pathfinding in 32x32 grid with unit collision radius
@@
- Target scoring: [0-10] based on type (colossus 9 > baneling 8 > zealot 7 > stalker 6 > hydralisk 5 > marauder 4 > marine 3 > zergling 2 > medivac 1)/health (0-0.6)/distance (0-0.3)/last attacked (0.1)
@@
- Default action: Move to center of map, parse region of interest, random choice
Args:
obs (str): Observation string containing game state
import math
# Parse observation
obs_data = parse_obs(obs)
# Get set of available actions
valid_actions = obs_data.available_actions
if 0 in valid_actions:
return 0
def score_target(unit):
@@
”””Enhanced target scoring with improved kiting and formation control”””
if unit.health <= 0:
return -1
score = 0
@@
# Refined unit type priorities with enhanced threat scaling
unit_priorities = {
colossus’: 35.0, # Further increased priority
stalker’: 30.0, # Enhanced anti-armor focus
zealot’: 45.0, # Higher melee threat recognition
marine’: 45.0, # Balanced damage dealer priority
marauder’: 35.0, # Anti-armor specialist
medivac’: 30.0, # Support unit priority
hydralisk’: 30.0, # High priority for their sustained DPS
zergling’: 35.0, # Medium priority as swarm units
baneling’: 45.0, # Critical priority due to splash damage
# Dynamic matchup priorities with improved counter weighting
unit_counters = {
colossus’: {’colossus’: 1.2, stalker’: 1.0, zealot’: 1.5},
stalker’: {’colossus’: 1.2, stalker’: 1.0, zealot’: 1.5},
zealot’: {’colossus’: 1.2, stalker’: 1.0, zealot’: 1.5},
marine’: {’marine’: 1.5, medivac’: 1.0, marauder’: 1.2},
marauder’: {’marine’: 1.5, medivac’: 1.0, marauder’: 1.2},
medivac’: {’marine’: 1.5, medivac’: 1.0, marauder’: 1.2},
hydralisk’: {’hydralisk’: 1.0, zergling’: 1.2, baneling’: 1.5},
zergling’: {’hydralisk’: 1.2, zergling’: 1.5, baneling’: 1.0},
baneling’: {’hydralisk’: 1.2, zergling’: 1.5, baneling’: 1.0},
base_priority = unit_priorities.get(unit.unit_type.lower(), 5.0)
is_ranged = unit.unit_type.lower() not in [’zealot’, zergling’, baneling’]
@@
own_is_ranged = obs_data.own_unit_type.lower() not in [’zealot’, zergling’, baneling’]
# Enhanced threat assessment with improved melee handling
matchup_mult = unit_counters.get(obs_data.own_unit_type.lower(), {}).get(unit.unit_type.lower(), 1.0)
base_priority *= matchup_mult
if hasattr(unit, can_attack’): # Enemy unit
score = base_priority
distance_factor = max((1 - unit.distance) + 1, 0.5)
score *= distance_factor
if not own_is_ranged:
@@
range_ally = [ally for ally in obs_data.allies if ally.unit_type.lower() not in [’zealot’, zergling’, baneling’]]
if range_ally:
ally_x = sum(ally.position[0] for ally in range_ally) / len(range_ally)
ally_y = sum(ally.position[1] for ally in range_ally) / len(range_ally)
ally_distance = ((ally_x - unit.position[0])**2 + (ally_y - unit.position[1])**2)**0.5
distance_factor = max((1 - ally_distance) + 1, 0.5)
score *= distance_factor
# Enhanced Position Analysis with improved spacing
position_x, position_y = unit.position
@@
def calculate_combat_power(units, radius=0.5): # Further reduced for tighter control
total_power = 0
ranged_count = 0
melee_count = 0
unit_positions = []
for u in units:
dist = ((u.position[0] - position_x)**2 +
(u.position[1] - position_y)**2)**0.5
unit_positions.append(u.position)
if dist <= radius:
base_power = unit_priorities.get(u.unit_type.lower(), 5.0)
# Unit type specific power calculation
if u.unit_type.lower() in [’zealot’, zergling’, baneling’]:
melee_count += 1
if melee_count >= 2:
base_power *= 1.4
else:
ranged_count += 1
base_power *= 1.3
# Health-based power scaling
health_factor = 1.5 if u.health > 0.7 else 1.0 if u.health > 0.4 else 0.6
position_factor = 1.3 - (dist/radius)
total_power += base_power * health_factor * position_factor
# Enhanced formation cohesion calculation
cohesion = 0
if len(unit_positions) > 2:
center_x = sum(p[0] for p in unit_positions) / len(unit_positions)
center_y = sum(p[1] for p in unit_positions) / len(unit_positions)
avg_dist = sum(((p[0] - center_x)**2 + (p[1] - center_y)**2)**0.5
for p in unit_positions) / len(unit_positions)
max_desired_dist = 0.3 # Tighter formation control
cohesion = 2.0 / (1.0 + (avg_dist / max_desired_dist))
return total_power * (1 + cohesion), melee_count, ranged_count
@@
ally_power, ally_swarms, ally_ranged = calculate_combat_power(obs_data.allies)
@@
enemy_power, enemy_swarms, enemy_ranged = calculate_combat_power(obs_data.enemies)
# Improved Focus Fire Logic with enhanced commitment
num_attackers = sum(1 for ally in obs_data.allies
if ally.last_action >= 6 and ally.last_action - 6 == unit.id)
if unit.id == obs_data.last_action - 6:
persistence_bonus = 2.0 # Stronger target commitment
score *= persistence_bonus
if num_attackers > 0:
focus_bonus = 1.2 ** num_attackers # Enhanced focus fire emphasis
# if num_attackers too high, discourage prevent overcommitment
@@
if num_attackers >= 3 and unit.id != obs_data.last_action - 6 and obs_data.own_unit_type.lower() in [’zergling’, baneling’]:
focus_bonus = 0.5
score *= focus_bonus
# Improved Combat Advantage Factor
advantage_factor = 1.0
if ally_power > enemy_power * 1.3:
advantage_factor = 1.2 # More aggressive advantage pursuit
if ally_swarms >= 3:
advantage_factor *= 1.2
# prioritize isolated enemies
if (enemy_swarms + enemy_ranged) == 1:
advantage_factor *= 2.0
elif (ally_swarms + ally_ranged) > (enemy_swarms + enemy_ranged):
advantage_factor *= 1.2
score *= advantage_factor
# Health factor
health_factor = (1 - unit.health) + 1
score *= health_factor
else: # Ally unit
score = base_priority
# Improved Support Priority
health_factor = (1 - unit.health) + 1
score *= health_factor
distance_factor = max((1 - unit.distance) + 1, 0.5)
score *= distance_factor
return score
def control_logic():
# Medivac units control logic
if obs_data.own_unit_type.lower() == medivac’:
attack_actions = [a for a in valid_actions if a >= 6]
# If there are allies
if obs_data.allies:
lowest_health_ally = min(obs_data.allies, key=lambda x: x.health)
# If there are both allies and enemies
if obs_data.enemies:
enemy_in_range = [enemy for enemy in obs_data.enemies if enemy.distance < 1]
@@
# Check if any melee ally, if so and last action is not attack, move to center of melee allies
@@
melee_ally = [ally for ally in obs_data.allies if ally.unit_type.lower() in [’zealot’, zergling’, baneling’]]
if melee_ally:
# Move to center of melee allies
ally_x = sum(ally.position[0] for ally in melee_ally) / len(melee_ally)
ally_y = sum(ally.position[1] for ally in melee_ally) / len(melee_ally)
else:
ally_x = sum(ally.position[0] for ally in obs_data.allies) / len(obs_data.allies)
ally_y = sum(ally.position[1] for ally in obs_data.allies) / len(obs_data.allies)
# Calculate retreat position
if len(enemy_in_range) > 0:
enemy_center = (sum(e.position[0] for e in enemy_in_range) / len(enemy_in_range),
sum(e.position[1] for e in enemy_in_range) / len(enemy_in_range))
else:
enemy_center = (sum(e.position[0] for e in obs_data.enemies) / len(obs_data.enemies),
sum(e.position[1] for e in obs_data.enemies) / len(obs_data.enemies))
dx = ally_x - enemy_center[0]
dy = ally_y - enemy_center[1]
distance = (dx ** 2 + dy ** 2) ** 0.5
safe_x = ally_x + (dx / abs(dx)) * 2/obs_data.own_sight_range if dx != 0 else ally_x
safe_y = ally_y + (dy / abs(dy)) * 2/obs_data.own_sight_range if dy != 0 else ally_y
distance = (safe_x ** 2 + safe_y ** 2) ** 0.5
criterion = 5/obs_data.own_sight_range
if obs_data.last_action >= 6 or len(attack_actions) == 0:
target_angle = math.atan2(enemy_center[1], enemy_center[0])
safe_angle = math.atan2(ally_y, ally_x)
angle_diff = abs(target_angle - safe_angle)
if distance > criterion and (math.pi/9 < angle_diff < 17*math.pi/9):
path_action = find_path(obs_data, safe_x, safe_y)
if path_action:
return path_action
target_scores = {ally.id: score_target(ally) for ally in obs_data.allies}
# Check if there are same max score targets
max_score = max(target_scores.values())
max_score_target_ids = [target_id for target_id, score in target_scores.items() if score == max_score]
max_score_targets = [ally for ally in obs_data.allies if ally.id in max_score_target_ids]
closest_ally = min(obs_data.allies, key=lambda x: x.distance)
if len(max_score_targets) > 1:
# Chose the closest target
best_target = min(max_score_targets, key=lambda x: x.distance)
else:
best_target = max_score_targets[0]
if (best_target.id + 6) in valid_actions and 0 < best_target.health < 0.9:
return heal(best_target.id)
elif (closest_ally.id + 6) in valid_actions and 0 < closest_ally.health < 0.9:
return heal(closest_ally.id)
elif (lowest_health_ally.id + 6) in valid_actions and 0 < lowest_health_ally.health < 0.9:
return heal(lowest_health_ally.id)
else:
# Move to the target
dx = best_target.position[0]
dy = best_target.position[1]
@@
path_action = find_path(obs_data, dx, dy, target_type=best_target.unit_type.lower())
if path_action:
return path_action
# If there are no allies
else:
# If there are only enemies
if obs_data.enemies:
enemy_x = sum(e.position[0] for e in obs_data.enemies) / len(obs_data.enemies)
enemy_y = sum(e.position[1] for e in obs_data.enemies) / len(obs_data.enemies)
target_x = - enemy_x
target_y = - enemy_y
g_x = target_x * obs_data.own_sight_range + (obs_data.own_position[0] * 32)
g_y = target_y * obs_data.own_sight_range + (obs_data.own_position[1] * 32)
if not (0 <= g_x <= 32 and 0 <= g_y <= 32):
target_x = (0.5 - obs_data.own_position[0]) * 32 / obs_data.own_sight_range
target_y = (0.5 - obs_data.own_position[1]) * 32 / obs_data.own_sight_range
path_action = find_path(obs_data, target_x, target_y)
if path_action:
return path_action
# Melee units control logic

### v2

@@ Appendix E Environment Settings and More Results
Appendix F Prompts used in COMPASS
@@ 5. Are there any ally units visible, either observation or minimap?
1. If your chosen region of interest is a unit, format the output as "[Enemy/Ally] #[target_id]" (e.g., "Enemy #0" for enemy unit with ID 0, "Ally #1" for ally unit with ID 1)
2. If your chosen region of interest is location, format the output as "Location: [direction]" where direction must be one of: "North", "Northeast", "East", "Southeast", "South", "Southwest", "West", "Northwest", "Center" (e.g., "Location: Northeast")
@@
Prompt for Task Reasoning You are an AI assistant helping with academic research in the StarCraft II’s SMAC (StarCraft Multi-Agent Challenge) environment, controlling a $<$unit_type$>$ unit with ID <$unit_{i}d$> in micromanagement scenarios $<$scenario_name$>$ to help your team defeat the enemy forces. You operate under decentralized execution with partial observability, making decisions based only on local observations within your unit’s field of view. You will be sequentially given $<$event_count$>$ screenshots and corresponding descriptions of recent events. You will also be given a summary of the history that happened before the last screenshot. By analyzing these inputs, you gain a comprehensive understanding of the current context and situation within the game. You should assist in summarizing the next immediate task to do in SMACv2. Your ultimate goal is to help your team defeat the enemy forces as quickly as possible.
@@
Based on the comprehensive game state analysis and team context, decompose the primary objective of "defeat all enemy units" into ONE specific tactical sub-task that enhances either target prioritization (score_target) or behavior control (control_logic). This sub-task should be concrete, implementable, and aligned with team coordination. Consider the following in your task decomposition:
@@ 2. Team Context:
"Adjust [scoring weight/multiplier/threshold] to [specific combat calculation] based on [unit composition + battle state] where [precise condition]"
@@
"Implement [unit movement pattern/formation/targeting] when [combat state + ally positions] satisfy [precise conditions]" Task Requirements:
@@
Prompt for Skill Generation You are an AI assistant helping with academic research in the StarCraft II’s SMAC (StarCraft Multi-Agent Challenge) environment, controlling a $<$unit_type$>$ unit with ID <$unit_{i}d$> in micromanagement scenarios $<$scenario_name$>$ to help your team defeat the enemy forces. You operate under decentralized execution with partial observability, making decisions based only on local observations within your unit’s field of view. Your task is to enhance combat effectiveness:
@@ 1. Analyze the provided script’s effectiveness
2. If the improved function is score_target(unit), there is exactly one parameter named "unit".
@@
Prompt for Actor You are an AI assistant helping with academic research in the StarCraft II’s SMAC (StarCraft Multi-Agent Challenge) environment, controlling a $<$unit_type$>$ unit with ID <$unit_{i}d$> in micromanagement scenarios $<$scenario_name$>$ to help your team defeat the enemy forces. You operate under decentralized execution with partial observability, making decisions based only on local observations within your unit’s field of view. Utilizing this insight, you are tasked with identifying the most suitable skill to take next, given the current task. You control the game unit and can execute skills from the available skill set. Upon evaluating the provided information, your role is to articulate the precise skill you would deploy, considering the game’s present circumstances, and specify any necessary parameters for implementing that skill:
@@
3. The ONLY required keyword parameter is "obs: str" - you MUST include this parameter as "obs=’current’" in every skill. The actual observation will be automatically injected at runtime.
@@
1def race_melee_ranged_medivac_navi_A_star_score_type_default_center(obs: str):
2 """
@@
3 Zealot/Zergling/Baneling/Colossus/Stalker/Hydralisk/Marauder/Marine/Medivac Controls Logic:
4 Medivac:
5 - Heals allies below 100% HP
6 - Maintains 0.75 sight range from enemies
7 - Centers between allies when no healing targets
8
9 Melee (Zealot/Zergling/Baneling):
10 - Attacks highest threat target within 0.7 sight range
11 - Pursues targets using A* pathfinding
12 - Groups with allies at >0.7 distance threshold
13
14 Ranged (Colossus/Stalker/Hydralisk/Marauder/Marine):
15 - Kites melee units at max_range - 0.05
16 - Focus fires targets shared with 2+ allies
17 - Maintains position behind melee allies
18
19 Key Implementation:
20 - Pathfinding: A* pathfinding in 32x32 grid with unit collision radius
@@ 19 Key Implementation:
21 - Target scoring: [0-10] based on type (colossus 9 > baneling 8 > zealot 7 > stalker 6 > hydralisk 5 > marauder 4 > marine 3 > zergling 2 > medivac 1)/health (0-0.6)/distance (0-0.3)/last attacked (0.1)
@@
22 - Default action: Move to center of map, parse region of interest, random choice
23
24 Args:
25 obs (str): Observation string containing game state
26 """
27 import math
28 # Parse observation
29 obs_data = parse_obs(obs)
30 # Get set of available actions
31 valid_actions = obs_data.available_actions
32
33 if 0 in valid_actions:
34 return 0
35
36 def score_target(unit):
@@ 24 Args:
37 """Enhanced target scoring with improved kiting and formation control"""
38 if unit.health <= 0:
39 return -1
40
41 score = 0
42
@@
43 # Refined unit type priorities with enhanced threat scaling
44 unit_priorities = {
45 ’colossus’: 35.0, # Further increased priority
46 ’stalker’: 30.0, # Enhanced anti-armor focus
47 ’zealot’: 45.0, # Higher melee threat recognition
48
49 ’marine’: 45.0, # Balanced damage dealer priority
50 ’marauder’: 35.0, # Anti-armor specialist
51 ’medivac’: 30.0, # Support unit priority
52
53 ’hydralisk’: 30.0, # High priority for their sustained DPS
54 ’zergling’: 35.0, # Medium priority as swarm units
55 ’baneling’: 45.0, # Critical priority due to splash damage
56 }
57
58 # Dynamic matchup priorities with improved counter weighting
59 unit_counters = {
60 ’colossus’: {’colossus’: 1.2, ’stalker’: 1.0, ’zealot’: 1.5},
61 ’stalker’: {’colossus’: 1.2, ’stalker’: 1.0, ’zealot’: 1.5},
62 ’zealot’: {’colossus’: 1.2, ’stalker’: 1.0, ’zealot’: 1.5},
63
64 ’marine’: {’marine’: 1.5, ’medivac’: 1.0, ’marauder’: 1.2},
65 ’marauder’: {’marine’: 1.5, ’medivac’: 1.0, ’marauder’: 1.2},
66 ’medivac’: {’marine’: 1.5, ’medivac’: 1.0, ’marauder’: 1.2},
67
68 ’hydralisk’: {’hydralisk’: 1.0, ’zergling’: 1.2, ’baneling’: 1.5},
69 ’zergling’: {’hydralisk’: 1.2, ’zergling’: 1.5, ’baneling’: 1.0},
70 ’baneling’: {’hydralisk’: 1.2, ’zergling’: 1.5, ’baneling’: 1.0},
71
72 }
73
74 base_priority = unit_priorities.get(unit.unit_type.lower(), 5.0)
75 is_ranged = unit.unit_type.lower() not in [’zealot’, ’zergling’, ’baneling’]
@@
76 own_is_ranged = obs_data.own_unit_type.lower() not in [’zealot’, ’zergling’, ’baneling’]
77
78 # Enhanced threat assessment with improved melee handling
79 matchup_mult = unit_counters.get(obs_data.own_unit_type.lower(), {}).get(unit.unit_type.lower(), 1.0)
80 base_priority *= matchup_mult
81
82 if hasattr(unit, ’can_attack’): # Enemy unit
83 score = base_priority
84
85 distance_factor = max((1 - unit.distance) + 1, 0.5)
86 score *= distance_factor
87
88 if not own_is_ranged:
@@
89 range_ally = [ally for ally in obs_data.allies if ally.unit_type.lower() not in [’zealot’, ’zergling’, ’baneling’]]
90 if range_ally:
91 ally_x = sum(ally.position[0] for ally in range_ally) / len(range_ally)
92 ally_y = sum(ally.position[1] for ally in range_ally) / len(range_ally)
93 ally_distance = ((ally_x - unit.position[0])**2 + (ally_y - unit.position[1])**2)**0.5
94 distance_factor = max((1 - ally_distance) + 1, 0.5)
95 score *= distance_factor
96
97 # Enhanced Position Analysis with improved spacing
98 position_x, position_y = unit.position
99
@@
100 def calculate_combat_power(units, radius=0.5): # Further reduced for tighter control
101 total_power = 0
102 ranged_count = 0
103 melee_count = 0
104 unit_positions = []
105
106 for u in units:
107 dist = ((u.position[0] - position_x)**2 +
108 (u.position[1] - position_y)**2)**0.5
109 unit_positions.append(u.position)
110
111 if dist <= radius:
112 base_power = unit_priorities.get(u.unit_type.lower(), 5.0)
113
114 # Unit type specific power calculation
115 if u.unit_type.lower() in [’zealot’, ’zergling’, ’baneling’]:
116 melee_count += 1
117 if melee_count >= 2:
118 base_power *= 1.4
119 else:
120 ranged_count += 1
121 base_power *= 1.3
122
123 # Health-based power scaling
124 health_factor = 1.5 if u.health > 0.7 else 1.0 if u.health > 0.4 else 0.6
125 position_factor = 1.3 - (dist/radius)
126
127 total_power += base_power * health_factor * position_factor
128
129 # Enhanced formation cohesion calculation
130 cohesion = 0
131 if len(unit_positions) > 2:
132 center_x = sum(p[0] for p in unit_positions) / len(unit_positions)
133 center_y = sum(p[1] for p in unit_positions) / len(unit_positions)
134 avg_dist = sum(((p[0] - center_x)**2 + (p[1] - center_y)**2)**0.5
135 for p in unit_positions) / len(unit_positions)
136 max_desired_dist = 0.3 # Tighter formation control
137 cohesion = 2.0 / (1.0 + (avg_dist / max_desired_dist))
138
139 return total_power * (1 + cohesion), melee_count, ranged_count
140
@@
141 ally_power, ally_swarms, ally_ranged = calculate_combat_power(obs_data.allies)
@@
142 enemy_power, enemy_swarms, enemy_ranged = calculate_combat_power(obs_data.enemies)
143
144 # Improved Focus Fire Logic with enhanced commitment
145 num_attackers = sum(1 for ally in obs_data.allies
146 if ally.last_action >= 6 and ally.last_action - 6 == unit.id)
147
148 if unit.id == obs_data.last_action - 6:
149 persistence_bonus = 2.0 # Stronger target commitment
150 score *= persistence_bonus
151
152 if num_attackers > 0:
153 focus_bonus = 1.2 ** num_attackers # Enhanced focus fire emphasis
154 # if num_attackers too high, discourage prevent overcommitment
@@
155 if num_attackers >= 3 and unit.id != obs_data.last_action - 6 and obs_data.own_unit_type.lower() in [’zergling’, ’baneling’]:
156 focus_bonus = 0.5
157 score *= focus_bonus
158
159 # Improved Combat Advantage Factor
160 advantage_factor = 1.0
161 if ally_power > enemy_power * 1.3:
162 advantage_factor = 1.2 # More aggressive advantage pursuit
163 if ally_swarms >= 3:
164 advantage_factor *= 1.2
165
166 # prioritize isolated enemies
167 if (enemy_swarms + enemy_ranged) == 1:
168 advantage_factor *= 2.0
169 elif (ally_swarms + ally_ranged) > (enemy_swarms + enemy_ranged):
170 advantage_factor *= 1.2
171
172 score *= advantage_factor
173
174 # Health factor
175 health_factor = (1 - unit.health) + 1
176 score *= health_factor
177
178 else: # Ally unit
179 score = base_priority
180
181 # Improved Support Priority
182 health_factor = (1 - unit.health) + 1
183 score *= health_factor
184
185 distance_factor = max((1 - unit.distance) + 1, 0.5)
186 score *= distance_factor
187
188 return score
189
190 def control_logic():
191 # Medivac units control logic
192 if obs_data.own_unit_type.lower() == ’medivac’:
193 attack_actions = [a for a in valid_actions if a >= 6]
194 # If there are allies
195 if obs_data.allies:
196 lowest_health_ally = min(obs_data.allies, key=lambda x: x.health)
197 # If there are both allies and enemies
198 if obs_data.enemies:
199 enemy_in_range = [enemy for enemy in obs_data.enemies if enemy.distance < 1]
@@
200 # Check if any melee ally, if so and last action is not attack, move to center of melee allies
@@
201 melee_ally = [ally for ally in obs_data.allies if ally.unit_type.lower() in [’zealot’, ’zergling’, ’baneling’]]
202 if melee_ally:
203 # Move to center of melee allies
204 ally_x = sum(ally.position[0] for ally in melee_ally) / len(melee_ally)
205 ally_y = sum(ally.position[1] for ally in melee_ally) / len(melee_ally)
206 else:
207 ally_x = sum(ally.position[0] for ally in obs_data.allies) / len(obs_data.allies)
208 ally_y = sum(ally.position[1] for ally in obs_data.allies) / len(obs_data.allies)
209 # Calculate retreat position
210 if len(enemy_in_range) > 0:
211 enemy_center = (sum(e.position[0] for e in enemy_in_range) / len(enemy_in_range),
212 sum(e.position[1] for e in enemy_in_range) / len(enemy_in_range))
213 else:
214 enemy_center = (sum(e.position[0] for e in obs_data.enemies) / len(obs_data.enemies),
215 sum(e.position[1] for e in obs_data.enemies) / len(obs_data.enemies))
216
217 dx = ally_x - enemy_center[0]
218 dy = ally_y - enemy_center[1]
219
220 distance = (dx ** 2 + dy ** 2) ** 0.5
221
222 safe_x = ally_x + (dx / abs(dx)) * 2/obs_data.own_sight_range if dx != 0 else ally_x
223 safe_y = ally_y + (dy / abs(dy)) * 2/obs_data.own_sight_range if dy != 0 else ally_y
224
225 distance = (safe_x ** 2 + safe_y ** 2) ** 0.5
226 criterion = 5/obs_data.own_sight_range
227 if obs_data.last_action >= 6 or len(attack_actions) == 0:
228 target_angle = math.atan2(enemy_center[1], enemy_center[0])
229 safe_angle = math.atan2(ally_y, ally_x)
230 angle_diff = abs(target_angle - safe_angle)
231 if distance > criterion and (math.pi/9 < angle_diff < 17*math.pi/9):
232 path_action = find_path(obs_data, safe_x, safe_y)
233 if path_action:
234 return path_action
235 target_scores = {ally.id: score_target(ally) for ally in obs_data.allies}
236 # Check if there are same max score targets
237 max_score = max(target_scores.values())
238 max_score_target_ids = [target_id for target_id, score in target_scores.items() if score == max_score]
239 max_score_targets = [ally for ally in obs_data.allies if ally.id in max_score_target_ids]
240 closest_ally = min(obs_data.allies, key=lambda x: x.distance)
241 if len(max_score_targets) > 1:
242 # Chose the closest target
243 best_target = min(max_score_targets, key=lambda x: x.distance)
244 else:
245 best_target = max_score_targets[0]
246 if (best_target.id + 6) in valid_actions and 0 < best_target.health < 0.9:
247 return heal(best_target.id)
248 elif (closest_ally.id + 6) in valid_actions and 0 < closest_ally.health < 0.9:
249 return heal(closest_ally.id)
250 elif (lowest_health_ally.id + 6) in valid_actions and 0 < lowest_health_ally.health < 0.9:
251 return heal(lowest_health_ally.id)
252 else:
253 # Move to the target
254 dx = best_target.position[0]
255 dy = best_target.position[1]
@@
256 path_action = find_path(obs_data, dx, dy, target_type=best_target.unit_type.lower())
257 if path_action:
258 return path_action
259 # If there are no allies
260 else:
261 # If there are only enemies
262 if obs_data.enemies:
263 enemy_x = sum(e.position[0] for e in obs_data.enemies) / len(obs_data.enemies)
264 enemy_y = sum(e.position[1] for e in obs_data.enemies) / len(obs_data.enemies)
265 target_x = - enemy_x
266 target_y = - enemy_y
267
268 g_x = target_x * obs_data.own_sight_range + (obs_data.own_position[0] * 32)
269 g_y = target_y * obs_data.own_sight_range + (obs_data.own_position[1] * 32)
270 if not (0 <= g_x <= 32 and 0 <= g_y <= 32):
271 target_x = (0.5 - obs_data.own_position[0]) * 32 / obs_data.own_sight_range
272 target_y = (0.5 - obs_data.own_position[1]) * 32 / obs_data.own_sight_range
273
274 path_action = find_path(obs_data, target_x, target_y)
275 if path_action:
276 return path_action
277 # Melee units control logic